### v3

Abstract
Creating group choreography from music has gained attention in cultural entertainment and virtual reality, aiming to coordinate visually cohesive and diverse group movements. Despite increasing interest, recent works face challenges in achieving aesthetically appealing choreography, primarily for two key issues: multi-dancer collision and single-dancer foot slide. To address these issues, we propose a Trajectory-Controllable Diffusion (TCDiff), a novel approach that harnesses non-overlapping trajectories to facilitate coherent dance movements. Specifically, to tackle dancer collisions, we introduce a Dance-Beat Navigator capable of generating trajectories for multiple dancers based on the music, complemented by a Distance-Consistency loss to maintain appropriate spacing among trajectories within a reasonable threshold. To mitigate foot sliding, we present a Footwork Adaptor that utilizes trajectory displacement from adjacent frames to enable flexible footwork, coupled with a Relative Forward-Kinematic loss to adjust the positioning of individual dancers’ root nodes and joints. Extensive experiments demonstrate that our method achieves state-of-the-art results. Project page is available at https://wanluzhu.github.io/TCDiffusion/.
Keywords:
Multi-Dander Generation Collision and Slide Trajectory-Controllable Diffusion Dance-beat navigator Footwork Adaptor
1 Introduction
Dance, renowned as one of the most expressive art forms, holds significant influence across cultural, cinematic, and academic domains [5, 56, 32, 62]. Creating dances demands meticulous manual editing processes, prompting the development of automated learning models for choreographing dances. In recent times, music-driven choreography, particularly in solo dancer scenes [51, 52, 36, 35, 34, 55], has garnered considerable attention. As immersive interactive experiences continue to evolve, multi-person scenarios have become a common necessity [62], leading to the emergence of music-driven group choreography characterized by cohesive and diverse group movements [48]. Despite being acknowledged [30] and explored in a few studies [29, 62, 50], these methods inevitably encounter two significant problems:
Figure 1: Visualizations of two significant problems: multi-dancer collision (the yellow box) and single-dancer foot slide (the red box). In contrast, our result (shown the right picture) yields superior visual aesthetics, devoid of collisions and foot slides.
Multi-dancer collision. In many choreography works [30, 29, 62, 55], movements and coordinates of a dancer are typically cascaded to form their model input. However, this results in an extremely unbalanced input structure, as movements often encompass over 100 dimensions, whereas coordinates are limited to just 3 dimensions. In group choreography, while the coordinates of dancers may vary significantly, their movements often exhibit similarity, posing a challenge for models attempting to differentiate between them. This similarity in movements, coupled with feature imbalance, can lead to confusion in distinguishing between individual dancers. We refer to this phenomenon as dancer ambiguity, highlighting the susceptibility of dancer-related information to confusion. GDanceR [30] believes that people with closer positions tend to have higher correlation, and utilizes cross-entity attention to pay more attention to adjacent dancers. However, we found that dancers’ positions are typically equidistant. For example, in a three-person dance formation shaped like a triangle, the distance between each dancer is almost equal[48]. This leads to models struggling to resolve dancer ambiguity based on distance. GCD [29] uses a local attention mask to restrict the model’s focus to individuals only. However, dancers exhibit similar movements, with only the coordinates showing significant differences. Without special attention towards trajectory, this similarity can still lead to dancer ambiguity. Hence, existing multi-dancer models result in individuals overlapping with each other, as shown in Figure 1.
Single-dancer foot slide. Foot sliding occurs due to the models’ weakness in modeling the correlation between the global trajectory and local rotations of other body parts [61]. In multi-person dances, the dancer ambiguity exacerbates the difficulty of modeling the correlation between footwork and displacement.
To address these issues, in this paper, we propose a two-stage method, called Trajectory-Controllable Diffusion (TCDiff), that first generates dancers’ coordinates and then produces dance movements. For multi-dancer collision, we present a Dance-Beat Navigator (DBN) designed to extract coordination with significant feature differences and fit them separately to avoid dancer ambiguity. Specifically, its pivotal component involves integrating a Distance-Consistency loss to regulate the distances between dancers, thereby preventing collisions. Additionally, we also introduce a straightforward yet powerful fusion projection that integrates information across dancers, enhancing the model’s capacity to distinguish between dancer features. For single-dancer foot slide, we present a Footwork Adaptor that derives footwork by analyzing the displacement between adjacent frames in the trajectory. In addition, we introduce a Relative Forward-Kinematic (RFK) loss to enhance the connection between the root node and the other joints of a single dancer, thereby reinforcing the root-motion relationship. In summary, our main contributions are:
  • We propose a Dance-Beat Navigator that can generate distinct dancer trajectories by exploring a Distance-Consistency loss to avoid dancer collision.
  • We present a Footwork Adaptor that utilizes the trajectory shift of adjacent frames to create corresponding footwork, incorporating a Relative Forward-Kinematic loss to enhance the correlation between the root node and dance motion, thus mitigating single-dancer foot sliding.
  • Building upon these components, we develop a novel two-stage multi-dancer generation framework, Trajectory-Controllable Diffusion, to produce high-quality dance movement. Experimental results demonstrate the superiority of our approach over existing methods.
2 Related Work
2.1 Music-driven Single-Dancer Generation
To generate Single-dancer motion based on music is challenging[24]. Early methods based on motion retrieval paradigm [28, 11, 33, 43] suffer from producing deformed actions issues. Recent works utilize large datasets [31, 21, 57, 36] to synthesize motions using a deep learning approach, including auto-regressive models[4, 1, 59, 51] and generative models[9, 55, 35, 61]. In recent years, Diffusion based models [55, 35, 18, 53] which achieve both high diversity and fidelity have demonstrated state-of-the-art performance. Nevertheless, they tend to produce freezing motions and abrupt transitions during long-term generations. To address this, LongDanceDiff [61] utilizes a conditional diffusion model framework [18, 53] akin to autoregressive processing, enhancing the diffusion’s ability to capture temporal information. In addition to generating natural dance movements, the controllability of dance actions is also an important yet underexplored research area. Current works provide temporal and spatial control like genre control [26], text control [13, 34], and joint control[55, 46]. However, achieving consistent and plausible group effects with a single-dancer model is challenging, as multi-person dance requires considering the correlations between different dancers. Single-dancer models also struggle to differentiate features among multiple dancers.
2.2 Multi-person Generation
To the best of our knowledge, only a handful of studies [62, 30, 29, 50] have focused on generating movements for multiple dancers. Among these, GDanceR [30] employs an encoder-decoder framework to auto-regressively generate movements for more than two individuals. GCD [29] is based on a diffusion framework and utilizes a contrastive encoder to generate results that are either diverse or consistent. Human motion forecasting is closely related to multi-dancer generation. To establish interactions between individuals, convolution networks are commonly used [6, 7, 8, 54]. To predict time series, Recurrent neural networks (RNNs) [12, 23, 41] and Transformers [44, 58] are commonly used. However, existing models lack structures for handling dancer ambiguity. In this work, we propose TCDiff, which utilizes Fusion Projection to enhance the ability to distinguish between different dancer characteristics, effectively reducing confusion.
2.3 Multi-agent Trajectory Prediction
Trajectory models aim to understand how agents move. Past methods heavily rely on hand-crafted rules for describing motions and interactions, including the Gaussian Process[25], and Markov Models[27]. However, these methods struggle with complex real-world scenarios. Recent deep learning approaches for temporal modeling are LSTM[2] and their variants. To model complex interactions attention-based methods and graph-based approaches have been developed, such as SGCN[49] and STAR[63]. Predicting trajectories is challenging due to the multi-modality issue [14, 16, 3], where the same input can lead to different outcomes. Consequently, current methods aim to learn a distribution rather than a single trajectory, employing generative models like GANs[16] and CVAEs[22] and DDPMs[15]. However, dancer trajectories are complicated by reliance on music, adding a new factor and even stationary dancers show intricate positional variations. In this work, TCDiff utilizes our proposed Dance-Beat Navigator, an auto-regressive model that can generate smooth, continuous, and non-overlapping trajectories for dancers.
3 Methodology
In this section, we introduce an effective method to generate synchronized group dance movements from a music clip. Our method consists of a Dance-Beat Navigator (DBN) and a Trajectory-Controllable Diffusion (TCDiff), as shown in Figure 2. Before showing them, we first present some backgrounds.
3.1 Background
Motion and Music Representations. We represent motion features of a dancer utilizing a SMPL[38] pose $\boldsymbol{d}\in\mathbb{R}^{144}$ extracted from a 24-joint SMPL model in 6D rotation format [67], along with 4-dimensional foot contact labels $\boldsymbol{f}\in\mathbb{R}^{4}$ and a 3-dimensional root node $\boldsymbol{p}\in\mathbb{R}^{3}$ for the positions of the dancer. This results in a motion representation $\boldsymbol{x}=[\boldsymbol{f},\boldsymbol{p},\boldsymbol{d}]\in\mathbb{R}^{151}$. Note that compared to 3D keypoint representations [66, 64, 40], the use of rotation format tends to yield better motion consistency, as observed in [51, 52]. For the music feature, we utilize Librosa [42] to extract a representation $\mathcal{M}\in\mathbb{R}^{35}$, comprising a 1-dimensional envelope, 20-dimensional MFCC, 12-dimensional chroma, along with 1-dimensional one-hot peaks and 1-dimensional one-hot beats.
Music-driven Group Choreography. Given music information =$\mathcal{M}=\left\{\boldsymbol{m}_{i}\right\}_{i=1}^{L}$, group choreography is to generate a corresponding group dance movement sequence $\boldsymbol{x}=\left\{\boldsymbol{x}^{i}\right\}_{i=1}^{L}$, where $\boldsymbol{x}^{i}=\left\{\boldsymbol{x}^{i,c}\right\}_{c=1}^{C}$, and $\boldsymbol{x}^{i,c}=[\boldsymbol{d}^{i,c},\boldsymbol{f}^{i,c},\boldsymbol{p}^{i,c}]$, $L$ is the music clip length and $C$ is the number of dancers.
Figure 2: Our framework consists of two main components: the Dance-Beat Navigator (DBN) and Trajectory-Controllable Diffusion (TCDiff). To address dancer ambiguity, initially, we employ DBN to model dancer positions, as dancers’ coordinates exhibit distinct differences and are less prone to confusion. Subsequently, TCDiff utilizes this result for conditional diffusion to generate corresponding dance movements. During this process, a fusion projection enhances group information before inputting it into the multi-dance transformer, while a footwork adaptor adjusts the final footwork.
Diffusion Model. We generate dance movements via a diffusion-based method [18, 53], which establishes a Markov noising process that gradually contaminates clean data $\boldsymbol{x}_{0}$ into standard Gaussian noise $\boldsymbol{x}_{T}\sim\mathcal{N}(0,\boldsymbol{I})$ through $T$ pollution steps. The corruption process is defined as the forward process:
$q\left(\boldsymbol{x}_{t}\mid\boldsymbol{x}_{t-1}\right)=\mathcal{N}\left(\sqrt{\alpha_{t}}\boldsymbol{x}_{t-1},\left(1-\alpha_{t}\right)I\right),$ (1)
where $\alpha_{t}\in(0,1)$ are pre-defined hyper-parameters. In music-driven choreography generation, the music $\mathcal{M}$ is frequently integrated as a conditioning factor [55]. The motion creation involves reversing the forward diffusion process by estimating $\widehat{\boldsymbol{x}}_{\theta}\left(\boldsymbol{x}_{t},t,\mathcal{M}\right)\approx\boldsymbol{x}$ with model parameters $\theta$ for all $t$. Thus, the basic objective [18] is defined as
$\mathcal{L}_{\text{simple }}=\mathbb{E}_{\boldsymbol{x},t}\left[\left\|\boldsymbol{x}-\widehat{\boldsymbol{x}}_{\theta}\left(\boldsymbol{x}_{t},t,\mathcal{M}\right)\right\|_{2}^{2}\right].$ (2)
In this paper, we extend the diffusion model by incorporating dancers’ trajectories to produce more realistic dance movements.
3.2 Dance-Beat Navigator
To mitigate dancer ambiguity resulting from similar movements among dancers, we introduce the Dance-Beat Navigator (DBN) module, which aims to suppress the interference of similar movements and prioritize the modeling of coordinates, as shown in the left of Figure 2. Given the music sequence =$\mathcal{M}=\left\{\boldsymbol{m}_{i}\right\}_{i=1}^{L}$, the DBN module operates as follows: starting with a seed trajectory $\boldsymbol{p}^{0}=\{\boldsymbol{p}^{0,1},\ldots,\boldsymbol{p}^{0,C}\}$, it takes the coordinates $\boldsymbol{p}^{i}=\{\boldsymbol{p}^{i,1},\ldots,\boldsymbol{p}^{i,C}\}$ and the music frame $\boldsymbol{m}_{i}$ at sequence step $i$ as inputs to a music encoder [26] and a trajectory encoder, along with an MLP decoder, to recursively generate the next dancer coordinates $\widehat{\boldsymbol{p}}^{i+1}=\{\widehat{\boldsymbol{p}}^{(i+1),1},\ldots,\widehat{\boldsymbol{p}}^{(i+1),C}\}$. The final trajectory $\widehat{\boldsymbol{p}}=\{\widehat{\boldsymbol{p}}^{1},\ldots,\widehat{\boldsymbol{p}}^{L}\}$ is obtained by applying a Savgol filter [45] to smooth the dancer coordinates. The MLP decoder is implemented as a simple multilayer perceptron. Next, we describe the trajectory encoder.
Trajectory Encoder. The Trajectory Encoder is tasked with extracting detailed features from both the input music and coordinate sequences, which are then fed into the MLP decoder to generate trajectory predictions. For feature pre-processing, the music sequence undergoes processing by a music-MLP, while the coordinate sequence is inputted into an LSTM [19] to extract temporal features. Additionally, instead of relying on absolute positional encoding [10, 58] or naive positional encoding [20, 37], we utilize identity encoding (IE) and temporal positional encoding (TPE) [44] to capture temporal information. Furthermore, we introduce a trajectory attention module comprising Casual Attention [58, 47], effectively directing the model’s focus to past information through masking, in conjunction with an MLP network. Its formulation is defined as:
$\operatorname{TrajectoryAttention}=\operatorname{MLP}\left(\underbrace{\operatorname{Softmax}\left(\boldsymbol{M}\boldsymbol{P}^{T}/\sqrt{d}+\boldsymbol{B}\right)\boldsymbol{P}\times mask}_{\text{CausualAttention}}\right),\vspace{-1.3mm}$ (3)
where $\boldsymbol{M}$, $\boldsymbol{P}$ are the processed music and position features using both the music-MLP and the LSTM, respectively. The $mask$ is the causal mask with $mask_{i,j}=-\infty\times 1(i>j)+1(i\leq j)$, where $1(\cdot)$ is the indicator function. $\boldsymbol{B}$ is the bias, and $d$ is a scaling factor to ensure the stability of the model’s training process. We replicate the trajectory attention module $M$ times in this context.
Loss of DBN. The objectives of the DBN module are determined by our observations that the ground truth typically exhibits characteristics of continuity and non-overlap. We found that relying solely on the reconstruction loss $\mathcal{L}_{\text{recon}}$ [65] is insufficient to guarantee this, therefore we have applied velocity loss $\mathcal{L}_{\text{v}}$ [55] and introduced Distance-Consistency loss $\mathcal{L}_{DistCon}$ as additional measures. To ensure continuity, we apply the velocity loss $\mathcal{L}_{\text{v}}$ to restrict position variation. To further approximate the ground truth for reducing the overlapping prediction, we introduce a Distance-Consistency loss $\mathcal{L}_{DistCon}$, formulated as:
$\mathcal{L}_{DistCon}=\frac{1}{C-1}\sum_{w=1}^{L}\binom{C}{2}_{ij}\left\|\left(\boldsymbol{p}^{w,i}-\boldsymbol{p}^{w,j}\right)-\left(\widehat{\boldsymbol{p}}^{w,i}-\widehat{\boldsymbol{p}}^{w,j}\right)\right\|_{2}^{2},$ (4)
which ensures that the spacing among dancers within an appropriate range. The overall loss $\mathcal{L}_{dbn}$ for DBN is $\mathcal{L}_{dbn}=\mathcal{L}_{recon}+\lambda_{v}\mathcal{L}_{v}+\lambda_{DistCon}\mathcal{L}_{DistCon}$.
3.3 Trajectory-Controllable Diffusion
After extracting trajectories from the DBN, we introduce TCDiff to generate dancer movements characterized by non-overlapping steps and enhanced grip, in accordance with the provided trajectories. TCDiff consists of a Fusion Projection, a Multi-Dancer Transformer, and linear layers for input and output. To ensure non-overlapping movements, we leverage disjoint trajectories obtained from the DBN for conditional denoising. To mitigate foot-slide, we present the Footwork Adaptor within the Multi-Dancer Transformer to adjust foot movements based on trajectory information, thereby reducing foot-slide occurrences. To fully exploit the known trajectory feature, we refrain from introducing noise to the provided positional information $\widehat{\boldsymbol{p}}$, opting instead for conditional motion denoising. Sinec the linear layers are simple single linear networks, the Fusion Projection and Multi-Dancer Transformer modules are described as follow.
Fusion Projection we propose is a simple yet effective solution that tackles dancer ambiguity challenges via its unique feature processing. Fusion Projection can be composed of a simple MLP structure, which achieves group feature extraction by converting individual-agent information into a group-agent format, as shown in Figure 3. Relevant evidence is provided in Section 4.3.
Multi-Dancer Transformer initially utilizes a Decoder block[55] to generate raw motion, which is then refined by our Footwork Adaptor to obtain adapted motion. We divide the human body representation into upper and lower parts, and by replacing the raw motion with the lower part of the adapted motion, we obtain the corrected footwork result. To be more specific, given frame-level music information =$\mathcal{M}=\{\boldsymbol{m}_{i}\}_{i}^{L}$, the diffusion final time step $T$, and the offset of the given trajectory $\mathcal{V}=\{\boldsymbol{v}_{i}\}_{i}^{L}$ (e.g., velocity) as conditions, we first utilize the Decoder Block module $D(\cdot)$ from [55] for feature processing to derive raw motion $\boldsymbol{\widehat{r}}$:
$\boldsymbol{\widehat{r}}=D(\boldsymbol{x}_{T},T,\mathcal{M}).$ (5)
Subsequently, we propose Footwork Adaptor module $FA(\cdot)$ for correction to obtain the adapted motion $\boldsymbol{\widehat{a}}$.
$\boldsymbol{\widehat{a}}=FA(\boldsymbol{\widehat{r}},T,\mathcal{M},\mathcal{V}).$ (6)
The Footwork Adaptor consists of a linear layer and a ConcatSquashLinear layer, as shown in Figure 2, which has been proven to be effective in various coordinate prediction domains [14, 39]. We split $\boldsymbol{\widehat{r}}$ and $\boldsymbol{\widehat{a}}$ into upper and lower body $\boldsymbol{\widehat{r}}=\{\boldsymbol{\widehat{r}}_{upper},\boldsymbol{\widehat{r}}_{lower}\}$, and $\boldsymbol{\widehat{a}}=\{\boldsymbol{\widehat{a}}_{upper},\boldsymbol{\widehat{a}}_{lower}\}$. The down part that is closely related to footwork is picked as the final generated result:
$\widehat{x}_{0}=\boldsymbol{\widehat{r}}_{upper}\oplus\boldsymbol{\widehat{a}}_{lower}$ (7)
Figure 3: Fusion Projection. The vanilla projection perceives a single dancer as an agent (Individual-Agent). We consider a group of dancers as a collective agent (Group-Agent), enhancing our capability to extract group features effectively.
Conditional Motion Denoising. To fully leverage the provided trajectory information, instead of adding noise to all information like the original diffusion does, we exclude adding noise to trajectory information. For denoising, we do not denoise $\widehat{\boldsymbol{p}}$ in $\boldsymbol{x}=[\boldsymbol{d},\boldsymbol{f},\widehat{\boldsymbol{p}}]$, but treat it as a condition for motion denoising. At each step $t$ in forward processing, we concatenate the noised dance pose information $\boldsymbol{d{}_{t}}$, the noised contact label $\boldsymbol{f}_{t}$ and the given position $\mathcal{\widehat{\boldsymbol{p}}}$ into one vector:
$\boldsymbol{x}_{t}=\mathcal{\boldsymbol{d{}_{t}}\oplus\boldsymbol{f{}_{t}}\oplus\widehat{\boldsymbol{p}}}$ (8)
Because the above data format is consistent with the original, we can still use simple loss in Eq. 2 for optimization. At this point, motion $\boldsymbol{d}$ and contact label $\boldsymbol{f}$ are denoised while coordinates $\widehat{\boldsymbol{p}}$ is reconstructed. This enhances the model’s learning and memory capabilities for coordinates while improving its ability to extract features from trajectories.
Loss of the TCDiff. To enable trajectory-conditional generation, we introduce the Relative Forward-Kinematic loss $\mathcal{L}_{\mathrm{RFK}}$, which adjusts the individual dancers’ relative distance between root nodes and other body joints. The $\mathcal{L}_{\mathrm{RFK}}$ is formulated as:
$\mathcal{L}_{\mathrm{RFK}}=\frac{1}{L}\sum_{i=1}^{L}\left\|\left(FK\left(\boldsymbol{d}\right)-FK\left(\boldsymbol{p}\right)\right)-(FK(\widehat{\boldsymbol{d}})-FK(\widehat{\boldsymbol{p}}))\right\|_{2}^{2}.$ (9)
Here, $FK(\cdot)$ is the forward kinematic function that calculates the positions of joints given the 6D rotation motion. We adopt joint velocity loss $\mathcal{L}_{\text{vel }}$ and the foot contact loss $\mathcal{L}_{\text{contact }}$ from [55]. The overall objective of our proposed TCDiff is $\mathcal{L}_{\mathrm{TCDiff}}=\lambda_{\text{simple}}\mathcal{L}_{\text{simple}}+\lambda_{\text{RFK}}\mathcal{L}_{\text{RFK }}+\lambda_{\text{vel }}\mathcal{L}_{\text{vel }}+\lambda_{\text{contact }}\mathcal{L}_{\text{contact }}.$
4 Experiments
Figure 4: Our generation results in varying numbers of dancers. Our approach achieves both diversity and coherency.
4.1 Experimental Settings
Implementation Details. For the Dance-beat Navigator, the $\lambda_{v}=\lambda_{DistCon}=2$, and the hidden size of all module layers is set to 64. The Trajectory transformer, which is stacked with $M=6$ transformer layers, is equipped with 8 heads of attention. Additionally, both the LSTM model and the Music-MLP consist of 3 layers each. The Final-MLP processes the information passed to it through 4 layers, utilizing LeakyReLU non-linearity as the activation function. The sequence length $L=120$, the hidden dimension is 512, with $N=8$ layers and 8 heads of attention. We apply a 3-layer MLP as a Fusion Projection, followed by ReLU activation at each layer. Additionally, we stack $W=3$ Concat Squash Linear with a hidden size of $d_{csl}=128$ and $d_{ctx}=512$. The entire framework was trained on 4 Nvidia 4090 GPUs for 3 days. We use a single 4090 GPU to train the Dance-Beat Navigator for 26 hours, utilizing batch sizes of 750, 400, 256, and 170 for 2, 3, 4, and 5 dancers, respectively. Similarly, the TCDiff model was trained on 4 NVIDIA 4090 GPUs for 1 day, employing batch sizes of 60, 53, 32, and 20, in that order.
Dataset. Our research builds upon the AIOZ-GDance dataset [30], an extensive repository of group dance performances comprising 16.7 hours of synchronized music and 3D multi-dancer motion data. This dataset encompasses a diverse array of over 4000 dancers, spanning 7 distinct dance styles and 16 music genres. We adhere to the dataset’s prescribed training and testing partition in our experimental framework.
Compared methods. To the best of our understanding, there’s a scarcity of multi-dancer models proficient in generating more than two dancers. In our evaluation, we juxtapose our methodology with the leading benchmarks in the field, namely GDanceR [30] and GCD [29]. GDanceR, introduced alongside the AIOZ-GDance dataset, employs an auto-regressive model. On the other hand, GCD, the subsequent advancement, demonstrates the highest qualitative performance to date. Additionally, we incorporate EDGE [55], the prevailing single-dancer model, to underscore the superiority of our approach.
Figure 5: Comparison with GCD [29] and GDanceR [30] in group. Our model produces fewer dancer overlaps and generates more natural footsteps during position exchange, whereas baselines either result in collisions or a lack of foot movements during the exchange process.
Metrics. We simultaneously employ metrics for multi-dancers and single-dancers to evaluate our model. For multi-dancer assessment [30], we employ Group Motion Realism (GMR), measuring feature similarity via Frechet Inception Distance. Group Motion Correlation (GMC) evaluates coherence through cross-correlation between generated dancers. Trajectory Intersection Frequency (TIF) assesses collision frequency among dancers.
For single-dancer evaluation, Frechet Inception Distance (FID) [36, 17] quantifies the similarity between individual dances and ground-truth dances. Generation Diversity (Div) [36, 21] appraises the variety of dance movements using kinetic features. Motion-Music Consistency (MMC) [36] evaluates how well generated dances synchronize with the music beat. Physical Foot Contact score (PFC) [55] indicates the physical plausibility of footwork by considering the correlation between the center of mass and foot velocity.
4.2 Comparison to the State of the Art
Figure 6: Dancer trajectories generated by our model, GdanceR [30] and GCD [29]. We can see that GdanceR produces overlapping trajectories, while GCD causes confusion among dancers, particularly highlighted in the red-box region where a purple dancer overlaps with a blue dancer. Our model reduces overlaps, showcasing the effectiveness of our Dance-Beat Navigator.
Table 1: Performance comparison with the state-of-the-art models GDanceR [30], GCD [29] and CoDancers [60]. Leveraging the advancements in single-dancer modeling from EDGE[55], we adapted it into a multi-dancer model by training it on the AIOZ-GDance dataset. Our approach attained superior results on the Group-dance Metric, concurrently mitigating foot slip occurrences. Red and Blue indicate the best and the second best result.
Method Group-dance Metric Single-dance Metric
[…]
EDGE [55] 70.26 62.24 0.42 33.24 10.11 0.26 3.02
GDanceR [30] 46.01 77.22 0.20 49.37 9.34 0.25 3.05
GCD [29] 31.95 80.55 0.16 37.70 9.64 0.26 2.53
CoDancers [60] 26.10 74.05 0.10 23.98 9.48 0.25 3.26
TCDiff (Ours) 13.86 81.98 0.13 37.47 15.10 0.25 0.51
Qualitative Visual Comparison. The performance of our model with varying numbers of dancers is depicted in Figure 4, showcasing its ability to generate aesthetically pleasing results across different group sizes. Figure 5 illustrates the comparison between our model and the baseline in dancer swapping. This comparison highlights that our model produces results with fewer overlaps and more realistic dancer movements, attributed to our improved handling of dancer ambiguity. Specifically, while the baseline model overly emphasizes the similarity of dancers’ movements during temporal prediction, it overlooks differences in positional coordinates, leading to confusion in their positions. Moreover, baselines are affected by dancer ambiguity, impeding their ability to correlate footwork actions with positional changes, thus making it challenging to generate corresponding footwork actions and resulting in unnatural movements. More visualization examples can be found on our website.
Table 2: Detailed comparison with SOTA Methods across various numbers of dancers. We observed that the interactions between dancers in duets manifest greater diversity, posing heightened challenges for learning. Nonetheless, our model demonstrates notable advantages across different group sizes.
Method Dancers # Group-dance Metric Single-dance Metric
[…]
GDanceR[30] 2 48.77 78.23 0.203 50.84 9.01 0.248
3 45.88 75.61 0.224 49.69 9.75 0.247
4 44.63 76.83 0.161 47.19 9.54 0.246
5 44.76 75.22 0.198 49.75 9.04 0.247
GCD[29] 2 30.10 79.36 0.125 37.28 9.78 0.260
3 33.26 79.78 0.163 36.24 9.87 0.260
4 34.23 82.83 0.172 38.07 9.45 0.251
5 30.22 80.22 0.185 39.24 9.45 0.260
CoDancers[60] 2 24.53 72.88 0.080 26.31 9.01 0.251
3 27.23 74.34 0.084 24.85 9.15 0.254
4 26.44 75.34 0.097 25.76 9.43 0.258
5 26.34 74.22 0.113 25.45 9.77 0.253
TCDiff (Ours) 2 15.77 81.92 0.121 41.26 16.20 0.263
3 10.97 81.51 0.123 48.00 19.28 0.253
4 13.44 81.70 0.149 23.32 10.89 0.253
5 15.36 82.77 0.109 37.31 14.01 0.236
Figure 7: User study based on four criteria: motion realism, music-motion correlation, formation realism, and harmony of dancers. Our model has garnered greater user favor, showcasing our superiority in aesthetic appeal.
Quantitative Results. Table 1 compares the results between our model and the state-of-the-art models. In Table 2, we offer a comprehensive comparison of multi-dancer models across various dancer counts. Our model demonstrates superior performance across all Group-dance Metrics and particularly stands out in Div and PFC within the Single-dance Metric. This suggests our model’s capability to produce cohesive multi-person dance sequences while enhancing diversity. While the Single-dancer model EDGE[55] maintains competitive performance in solo dance metrics, it underperforms in multi-person dance metrics, as indicated by its high TIF value, suggesting a tendency for overlapping results due to confusion between dancer information, resulting in random swaps and overlaps. Conversely, the auto-regressive model GDanceR[30] tends to generate effects lacking movement, predominantly executing low-quality dance actions in a stationary manner, consistent with findings in [29]. Its high TIF value reflects challenges in handling dancer ambiguity. GCD [29] also confronts the challenge of dancer ambiguity, leading to unnatural transitions between movements, thereby contributing to its high TIF. Our approach addresses this issue by decoupling the generation of dancer coordinates and movements into two distinct stages, thereby mitigating dancer ambiguity and yielding improved outcomes. The findings from our user study, illustrated in Figure 7, further validate the visual superiority of our model.
Table 3: Ablation study of Footwork Adaptor (FA) and Fusion Projection (FP).
Method Group-dance Metric Single-dance Metric
[…]
w/o FA and FP 25.57 74.80 0.149 27.18 17.86 0.19 4.29
w/o FA 21.39 80.13 0.149 29.21 12.79 0.22 3.25
w/o FP 24.25 82.63 0.148 23.77 16.00 0.21 0.72
Full 13.86 81.98 0.126 37.47 15.10 0.25 0.51
Figure 8: Ablation study compares EDGE with and without Fusion Projection module. We offer dancers’ coordinates from time 0s to now, aiding in observing multi-dancer crossover. Without Fusion Projection, EDGE shows frequent dancer overlaps and swaps (shown by the purple box), suggesting dancer ambiguity. Fusion Projection improves the model’s ability to distinguish between dancers.
4.3 Ablation Study
Effectiveness of Fusion Projection. As depicted in Table 3, Fusion Projection results in noticeable improvements across various metrics such as GMR, GMC, MMC, and PFC. These enhancements signify the generation of more lifelike group dance sequences with better synchronization to the music beats. Note that the TIF value, primarily influenced by the Dance-Beat Navigator’s generated dancer positions, remains consistent across models.
Furthermore, our investigation reveals a substantial improvement in the model’s capacity to manage dancer ambiguity through Fusion Projection. Figure 8 illustrates the generation outcomes of EDGE[55] both with and without our module. Given that EDGE is originally designed for single-dancer generation, its inherent weakness in handling dancer ambiguity is evident. However, the integration of our Fusion Projection module notably fortifies its capability in addressing this challenge, thereby validating the efficacy of our approach.
Effectiveness of Footwork Adaptor. We highlight the importance of natural footwork in enhancing visual aesthetics [55]. Analysis from Table 3 reveals that our Footwork Adaptor markedly enhances the model’s PFC value. Moreover, when utilized independently, it boosts the model’s GMC, underscoring the Footwork Adaptor’s dual benefits: preventing foot slipping and augmenting dance coherence.
4.4 Application
Formation Control. Group formation plays a pivotal role in group dancing, with a carefully orchestrated arrangement elevating visual aesthetics significantly [48]. A controlled dance formation is crucial for meeting practical application requirements effectively. Our approach enables the generation of dance movements that precisely adhere to specified dancer positions. Users can input coordinates for dance generation, ensuring meticulous control over group spatial dynamics. Throughout the diffusion sampling process, we assign noise-infused motion data specific positional values, maintaining consistency between input and output trajectories and thereby ensuring controllability.
Figure 9: Position control results: Our model leverages user-provided dancer positions to facilitate position swaps, yielding sensible and manageable outcomes while also generating corresponding footwork movements—a capability beyond the existing models.
Position Exchange Control. In group dancing, positional exchange stands out as a distinctive behavior compared to solo dancing, as it involves multiple dancers seamlessly swapping positions while upholding the formation. This exchange contributes to a richer visual experience and adds novelty for the audience. Our model excels in leveraging user-provided trajectories for positional swaps, yielding both reasonable and controllable outcomes, and synchronizing them with corresponding footwork movements, as illustrated in Figure 9, a feat that existing models fail to accomplish.
5 Limitation
Our method faces certain limitations. Firstly, processing multi-dancer features extends training times and demands significant computational resources. Secondly, while our model excels in achieving strict coordination controllability, it could benefit from introducing variations based on the scene rather than adhering strictly to predetermined paths. Additionally, although the Dance-Beat Navigator effectively reduces collisions between dancers, solely learning individual dancer coordination without considering interactions with other joints exacerbates the multi-modality trajectory prediction issue. Enhancements in the performance of the Dance-Beat Navigator could significantly enrich the variety of movements in our dance model.
6 Conclusion
In this paper, we introduce TCDiff, an innovative framework designed to produce high-quality multi-dancer movement. Through extensive experimentation utilizing the AIOZ-GDance dataset, we showcase the superior performance of our model in both individual and group dance scenarios, establishing its position at the forefront of current research. Notably, our framework offers spatial control, empowering users to precisely dictate dancer coordinates. This feature facilitates Formation Control and Position Exchange Control, essential elements for generating group dances with remarkable precision and fluidity.